OpenStudy (anonymous):

A collection of quarters and dimes is worth $18.30. There are 111 coins in all. How many of each coin are there? _____Quarters _____Dimes @igreen

OpenStudy (anonymous):

72 quorters 3 dimes

OpenStudy (igreen):

We can write two equations and solve. 0.25q + 0.10d = 18.30 q + d = 111 Where q = quarters and d = dimes.

OpenStudy (igreen):

Here's the graph..what's the solution? @literally.kaylaa

OpenStudy (igreen):

@Algebruhh Don't give out answers, plus your answer is wrong. 72 + 3 = 75 coins total..it says there are 111 coins total.

OpenStudy (igreen):

The solution to the system of equation is (48, 63). (quarters, dimes) So how many quarters and dimes are there? @literally.kaylaa

OpenStudy (anonymous):

you two are making it too hard. 4 quarters equal a dollar. So you take 18 x 4 = 72. Then 3 dimes makes up the extra 30 c

undeadknight26 (undeadknight26):

^

OpenStudy (anonymous):

oh sorry I was working it out >.< @igreen

OpenStudy (anonymous):

oh, i didnt pay attention to how many coins you had to use. But people use this website for answers not help.

OpenStudy (igreen):

@Algebruhh 72 + 3 is 75..not 111. It says there needs to be a total of 111 coins. @Algebruhh @undeadknight26

undeadknight26 (undeadknight26):

I didnt say anything o.o just backin your forumla up o.o

OpenStudy (anonymous):

48 quarters. 63 dimes. Suck It !!!

OpenStudy (igreen):

Oh..well it looks like you posted and pointed at his comment.

OpenStudy (igreen):

@Algebruhh People are SUPPOSED to use this so people can GUIDE them to the question, not so they can have answers.